# A Decade of Systems for Human Data Interaction

### v1

@@ 1 Introduction
Human–Data Interaction (HDI) studies how people access, understand, manipulate, and manage data. It spans scenarios such as querying databases through natural language interfaces [li2014constructing], performing large-scale visual analytics [heer2010zoo], and managing personal or organizational data [mortier2014human]. Among these settings, data visualization interfaces are especially demanding: they must communicate complex information effectively and support expressive, low-latency interactions at the pace of human thought. As a result, visualization interfaces push the limits of both data systems and user interface design.
@@
Visualization, in contrast, starts with the user. As Card et al. write, it focuses on “interactive, visual representations of data to amplify cognition [card1999readings],” while Heer and Kandel emphasize the need to support exploration at the rate of human thought [heer2012interactive].” These definitions highlight how visual representation empowers users, but remain agnostic to critical system concerns—such as how data is modeled, how analyses are executed, or how the system is architected and abstracted—which are central to data management research.
Historically, the separation was effective. Visualization tools focus on user experience, while DBMSes optimize performance and scalability. Small datasets can be run in the browser [heer2006software, bostock2011d3, satyanarayan2017vega], while larger workloads simply send SQL to a backend and rely on the DBMS’s performance. Tools like Tableau and Looker embody this loose coupling, where the client translates user interactions into queries sent to backend data systems.
@@
This design dependency (Figure 1) goes beyond performance because new system abstractions can enable new classes of interaction design, while new visual designs or interaction requirements motivate new system designs. For example, streaming databases like Materialize enable real-time maintenance of predefined views but, in turn, constrain dashboards to interactions that efficiently map to those maintained views [mcsherry2013differential]. In short, system design shapes interaction design, and interaction design shapes system design .
@@ 2.1 Black-box Query Workload
A large body of prior work treats the interface as a black box that produces a stream of queries. This model is attractive because it mirrors how most visualization systems operate today: user interactions (e.g., brushing, filtering, or zooming) are translated into SQL queries that the underlying DBMS executes to update visible views. Under this model, the interface reduces to a traditional analytical workload, and decades of database research that reduce query latency and improve throughput should, in principle, directly benefit interactivity [^1] For a comprehensive review, see Battle et al. [battle2020structured].. This model also makes it straightforward to create interaction benchmarks [eichmann2020idebench, battle20benchmark].
Much of the systems literature on visualization performance adopts this viewpoint. For example, data cubes reduce the cost of common aggregation queries by pre-computing summaries over low-dimensional group-bys that are often small enough to cache in memory [liu2013immens, lins2013nanocubes, pahins2016hashedcubes]. Similarly, indexes accelerate pan-and-zoom interactions and nearest neighbor look ups in map-based or spatial interfaces [kyrix]. Because these structures align with familiar analytical patterns, they can be implemented directly on existing DBMSs or embedded engines. Mosaic exemplifies this approach by building partial data cubes via SQL and evaluating them client-side using an embedded DBMS [mosaic] rather than relying on bespoke data structures.
Unfortunately, simply reducing query latency does not guarantee better interactivity. Standard metrics such as p95 latency often fail to correlate with perceived responsiveness [jiang2018evaluating, rahman2020evaluating], and reflect a mismatch between system-level performance indicators and user experience. Treating the interface purely as a query workload source abstracts away critical aspects of interaction design, perception, and expectation. For example, a nominally simple action—like adjusting a slider—may trigger a burst of concurrent queries, many of which update off-screen or low-priority views. Executing all queries at once forces them to compete for shared resources and can delay the updates the user is actively attending to.
@@ 2.2 Interface Properties
As the limitations of the black-box workload model suggest, optimizing standard workload metrics does not always lead to fluid user experiences [jiang2018evaluating]. Modern data interfaces can easily generate enough queries to saturate compute, memory, and/or network resources. Thus, the challenge is not to simply execute every query quickly, but to allocate resources to the queries that best meet the user’s current needs.
@@ 2.2.1 Network Communication
Many interactive data interfaces are deployed as web-based client–server applications, where network latency and bandwidth often dominate response time. Prior work exploits this property to optimize the client-server communication. Classic approaches such as prefetching predict what the user is likely to do next and proactively send responses before the user’s request is issued. This is effective when requests are sparse and predictable, such as in document retrieval or web search, where task models and behavioral signals provide reliable cues and there is enough time between the prediction and user’s request to perform the prefetch [battle2016dynamic].
@@
Motivated by these constraints, our Khameleon [mohammed2020continuous] system reframes the problem: it guarantees immediate responses and treats fidelity as the optimization objective. Khameleon progressively encodes results so that any prefix can be rendered, giving the client flexibility to hedge across many possible future interactions or to store full-fidelity results for a few likely ones. Client messages (including explicit prefetch requests) serve only as hints that help the server allocate its network budget more effectively; the server proactively pushes partial results and remains decoupled from the client.
@@ 2.2.2 Approximation
The second property is to leverage imperfect user senses. Decades of work in perception and psychophysics, such as Weber’s just-noticeable difference (JND) [jnd, fechner1860elemente], seek to quantify the extent humans can detect numerical differences. These insights motivated the rich literature on approximate query processing (AQP) [aqp, agarwal2013blinkdb, hellerstein1997online], and more recently, on approximations targeted specifically at visualization.
Visualization-oriented approximation typically seeks to preserve visual rather than statistical fidelity. Because the final visualization has finite pixel resolution, techniques such as M4 [m4] and our work on OM3 [om3] exploit “resolution push-down” to pre-aggregate statistics at the pixel level so that the visualization can be reconstructed exactly from coarse summaries. This shifts the cost from network transfer to database computation. OM3 further provides progressive prefixes that render increasingly refined versions of the chart, mirroring Khameleon’s progressive encoding strategy.
When approximations diverge from exact results, systems must assume the types of errors users find acceptable and find ways to convey the uncertainty. However, users struggle to interpret statistical uncertainty [Hullman2019Why, Padilla2021Uncertainty, Kahneman1982Judgment]. Further, accuracy guarantees over individual aggregates are not enough, because they may not preserve higher-level patterns that guide visual interpretation. BlinkDB [agarwal2013blinkdb] addresses these issues by exposing user-specified error or time budgets, but users find it hard to set these parameters. Distribution-oriented methods seek to preserve shapes or density profiles [sampleseek, macke2021rapid], or to preserve properties of functions over the resulting visualization; for instance, Rahman et al. [rahman2017ve] preserves pairwise orderings in bar charts.
Unfortunately, perceptual error tolerances vary across people, tasks, and encodings and are not yet fully characterized. This makes property-driven approximations powerful but brittle. An open question is how to define perceptual models that are both general enough to apply across visualizations and precise enough to support optimization. Such models would allow systems to reason about approximation error in terms of perceived fidelity rather than statistical metrics. For instance, our work on perceptual functions (p-funks) [alabi2016pfunk] represents an early attempt to formalize these constraints. It specifies a just-noticeable difference (JND) model a function of mark properties (e.g., pixel height) so the system can guarantee that online approximations remain within perceptual bounds. This illustrates how interface-aware perceptual models can guide system behavior, but also highlights the need for more principled, general representations of visualization semantics.
@@ 2.2.3 A Deluge of Special Cases
Interface designers face a difficult and increasingly untenable position. To achieve interactive performance, they must choose layouts, encodings, and interaction patterns that align with a sprawling ecosystem of bespoke system optimizations—each tailored to a narrow performance sweet spot under specific interface or user assumptions. Simply understanding the subtle distinctions between dozens of approximation, caching, and scheduling strategies is challenging. Further deciding which combinations apply to a particular interface is beyond the reach of move developers.
@@ 2.3 Formal Representations & Design Independence
The brittle nature of special-case optimizations highlights the need for Design Independence: a principled way to represent interface semantics so the system, rather than the designer, can decide how to meet interactivity constraints. Design Independence would let interface designers rapidly iterate without being bottlenecked by checking with developers or architects about implementation feasibility [Walny2019DataCE], much as logical plans let query optimizers reason about execution without exposing physical details.
@@ 2.3.1 Physical Visualization Design
Physical Visualization Design (PVD) [chen2025physical] is our approach towards Design Independence. Similar to physical database design (PDD) [christian2011autoadmin, zilio2004db2, finkelstein1988physical], which selects physical structures and layouts to accelerate a representative query workload, PVD selects data structures, caching strategies, and execution plans to meet the latency constraints of a visualization interface. The key idea is to provide a formal yet lightweight representation of the interface that captures the data flow, interactions, and latency requirements. This representation is input to an automated optimizer can analyze to synthesize a feasible client–server–DBMS execution strategy and instantiate the architecture.
@@
At first glance, PVD should not be possible. PDD is already considered intractable due to challenges of cost estimation and a combinatorial search space, and solutions only attempt to reduce expected workload latency. PVD appears to be strictly harder because it guarantees wall-clock latency for interactions. Surprisingly, we find that these millisecond-to-second latency constraints actually simplify the search because they eliminate the hardest cases in query optimization: joins with large or unpredictable fan-out. If a join cannot guarantee bounded output size [armbrust2009scads], it cannot satisfy the latency constraints and is pruned. This dramatically reduces the search space, making PVD feasible despite its stronger guarantees.
@@ 2.3.2 An Instance of Physical Visualization Design
Our system, Jade [chen2025physical], is the first instantiation of PVD that focuses on choosing and placing data structures for interactive visualization interfaces. Because modern visualizations rely on bespoke data structures (e.g., cubes, sketches, indexes), Jade is designed to support an extensible library of custom data structures.
@@
Figure 3(d) uses Cube.build() to compute a byte array that encodes a dense cube, the Cache() operator to reuse the cube when \$a or \$b change, and Cube.eval(2001, 2020) uses the cube to compute the vote distribution between 2001 and 2002. Data structures also estimate their cost and size based on statistics, and specify a match() method to identify subplans they can replace.
@@ 3.1 From Query Workloads to New Interactions
While popular, this library largely reflects what today’s systems can execute efficiently, rather than the full range of analytical operations users perform. For example, what if the user wishes to compare data slices, e.g., “the average cost over time for 75-year-olds against the overall population?” Such comparisons are commonplace in OLAP and cube-based analytics [gray1997data, sarawagi2000i3] but have no counterpart in modern visualization systems because they require reasoning about the semantics and schemas of multiple derived views, not just editing a predicate in a single query.
@@
This gap motivated our View Composition Algebra work (VCA) [wu2022view], which treats views as first-class algebraic objects that can be composed, joined, merged, or differenced. Since views are the output of query pipelines, VCA operates over entire data-processing pipelines; this is in contrast to traditional interactions that modify SQL literals. Figure 8 demonstrates VCA in practice: the user drags the 2016 election results onto the 2020 results to compute their difference and visualize the result as a map—with no manual data wrangling. Figure 7 shows 4 illustrative VCA operators: (a) element-wise subtracting one marks in one view with its match in the other, (b) unioning the records in two views together, (c) computing the average of each date’s delays across all cities, and (d) implicitly joining the underlying tables to subtract the blue bar from every date.
Ideally, a user can freely compare any combination of data visualized in an interface. However, supporting such ad hoc comparisons in visual analysis tools is tricky because the specific data that will be visualized is often not known ahead of time and not all data is directly comparable. Thus, ad hoc comparisons require the system to dynamically assess whether a combination of views is valid to compose. From a data management perspective, this reduces to classic data integration problems: Are two views comparable? That is, do their schemas admit a semantically valid mapping (e.g., cost $\sim$ price)? If so, which tuples should be matched? This is entity resolution. Data integration is remains an open problem, and thus, advances in schema matching, type inference, and entity resolution directly expand what comparison interactions can support.
Comparison is just one example of interaction techniques made possible by data-management concepts. Another is explanatory interactions, which surface predicates that explain surprising visual patterns [wu2013scorpion]. But enabling explanations inside interactive systems required nearly a decade of systems innovations from our lab—to support fast lineage capture [Psallidas2018SmokeFL, Psallidas2018ProvenanceFI, Psallidas2018DemonstrationOS] and scalable lineage evaluation [mohammed2025fade]—that combines with prior work on fast explanation discovery algorithms [abuzaid2018diff, roy2015explaining, gathani2020debugging]. These systems-level innovations need to be in place before such interactions are feasible at the latency expected from modern interfaces.
@@ 3.2 From Query Workloads to New Interfaces
Our second project shows how the compositional structure in query workloads be used to support automatic interface synthesis to go beyond today’s template-based approaches to interface design.
@@
We formalized this idea as the Data Interface Grammar (DIG) [dig]. DIG models interfaces as grammars whose non-terminals correspond to interface choices and whose productions encode schema-aware SQL structure. DIG extends Parsing Expression Grammars (PEGs) [peg] with SQL- and schema-specific annotations that constrain allowable expressions. A valid interface is a visual representation that expresses this grammar—user interactions resolve non-terminals to produce a concrete parse tree, which corresponds to a valid query.
@@
This representation offers several benefits. First, applications don’t assemble raw SQL strings and every query conforms to the grammar. Second, grammar rewrites offer a simple mechanism to refactor and explore alternative interface designs. Third, DIG aligns naturally with existing data engineering practices. For instance, DBT pipelines already structure analytics as DAGs of SQL views, which DIG can express and directly map to interactive dashboards. Fourth, DIG shares structural similarities with the plan-based representations used in PVD (Section 2.3), making it amenable to automatic optimization. Finally, standard techniques such as grammar induction can infer DIG grammars from query traces, enabling interface synthesis—an approach we explored in a line of work called Precision Interfaces [pi, pi2, pinotebook, nl2interface].
DIG is also useful when using large language models (LLMs). Natural language queries often contain unresolved ambiguities in intent, schema interpretation, or operator choice [nlambig]. Rather than forcing the system to commit to a single query, the LLM can simply represent ambiguity as non-terminals and return an interface whose interactions navigate the space of ambiguities. This idea underlies NL2Interface [nl2interface], where LLMs generate a grammar expressing possible interpretations, and the user resolves them through interaction.

### v2

@@ 1 Introduction
Human–Data Interaction (HDI) studies how people access, understand, manipulate, and manage data. It spans scenarios such as querying databases through natural language interfaces [46], performing large-scale visual analytics [31], and managing personal or organizational data [54]. Among these settings, data visualization interfaces are especially demanding: they must communicate complex information effectively and support expressive, low-latency interactions at the pace of human thought. As a result, visualization interfaces push the limits of both data systems and user interface design.
@@
Visualization, in contrast, starts with the user. As Card et al. write, it focuses on “interactive, visual representations of data to amplify cognition [11],” while Heer and Kandel emphasize the need to support exploration at the rate of human thought [32].” These definitions highlight how visual representation empowers users, but remain agnostic to critical system concerns—such as how data is modeled, how analyses are executed, or how the system is architected and abstracted—which are central to data management research.
Historically, the separation was effective. Visualization tools focus on user experience, while DBMSes optimize performance and scalability. Small datasets can be run in the browser [30, 10, 67], while larger workloads simply send SQL to a backend and rely on the DBMS’s performance. Tools like Tableau and Looker embody this loose coupling, where the client translates user interactions into queries sent to backend data systems.
@@
This design dependency (Figure 1) goes beyond performance because new system abstractions can enable new classes of interaction design, while new visual designs or interaction requirements motivate new system designs. For example, streaming databases like Materialize enable real-time maintenance of predefined views but, in turn, constrain dashboards to interactions that efficiently map to those maintained views [51]. In short, system design shapes interaction design, and interaction design shapes system design .
@@ 2.1 Black-box Query Workload
A large body of prior work treats the interface as a black box that produces a stream of queries. This model is attractive because it mirrors how most visualization systems operate today: user interactions (e.g., brushing, filtering, or zooming) are translated into SQL queries that the underlying DBMS executes to update visible views. Under this model, the interface reduces to a traditional analytical workload, and decades of database research that reduce query latency and improve throughput should, in principle, directly benefit interactivity [^1] For a comprehensive review, see Battle et al. [9].. This model also makes it straightforward to create interaction benchmarks [22, 8].
Much of the systems literature on visualization performance adopts this viewpoint. For example, data cubes reduce the cost of common aggregation queries by pre-computing summaries over low-dimensional group-bys that are often small enough to cache in memory [49, 48, 56]. Similarly, indexes accelerate pan-and-zoom interactions and nearest neighbor look ups in map-based or spatial interfaces [75]. Because these structures align with familiar analytical patterns, they can be implemented directly on existing DBMSs or embedded engines. Mosaic exemplifies this approach by building partial data cubes via SQL and evaluating them client-side using an embedded DBMS [33] rather than relying on bespoke data structures.
Unfortunately, simply reducing query latency does not guarantee better interactivity. Standard metrics such as p95 latency often fail to correlate with perceived responsiveness [40, 62], and reflect a mismatch between system-level performance indicators and user experience. Treating the interface purely as a query workload source abstracts away critical aspects of interaction design, perception, and expectation. For example, a nominally simple action—like adjusting a slider—may trigger a burst of concurrent queries, many of which update off-screen or low-priority views. Executing all queries at once forces them to compete for shared resources and can delay the updates the user is actively attending to.
@@ 2.2 Interface Properties
As the limitations of the black-box workload model suggest, optimizing standard workload metrics does not always lead to fluid user experiences [40]. Modern data interfaces can easily generate enough queries to saturate compute, memory, and/or network resources. Thus, the challenge is not to simply execute every query quickly, but to allocate resources to the queries that best meet the user’s current needs.
@@ 2.2.1 Network Communication
Many interactive data interfaces are deployed as web-based client–server applications, where network latency and bandwidth often dominate response time. Prior work exploits this property to optimize the client-server communication. Classic approaches such as prefetching predict what the user is likely to do next and proactively send responses before the user’s request is issued. This is effective when requests are sparse and predictable, such as in document retrieval or web search, where task models and behavioral signals provide reliable cues and there is enough time between the prediction and user’s request to perform the prefetch [7].
@@
Motivated by these constraints, our Khameleon [53] system reframes the problem: it guarantees immediate responses and treats fidelity as the optimization objective. Khameleon progressively encodes results so that any prefix can be rendered, giving the client flexibility to hedge across many possible future interactions or to store full-fidelity results for a few likely ones. Client messages (including explicit prefetch requests) serve only as hints that help the server allocate its network budget more effectively; the server proactively pushes partial results and remains decoupled from the client.
@@ 2.2.2 Approximation
The second property is to leverage imperfect user senses. Decades of work in perception and psychophysics, such as Weber’s just-noticeable difference (JND) [80, 24], seek to quantify the extent humans can detect numerical differences. These insights motivated the rich literature on approximate query processing (AQP) [27, 2, 34], and more recently, on approximations targeted specifically at visualization.
Visualization-oriented approximation typically seeks to preserve visual rather than statistical fidelity. Because the final visualization has finite pixel resolution, techniques such as M4 [41] and our work on OM3 [79] exploit “resolution push-down” to pre-aggregate statistics at the pixel level so that the visualization can be reconstructed exactly from coarse summaries. This shifts the cost from network transfer to database computation. OM3 further provides progressive prefixes that render increasingly refined versions of the chart, mirroring Khameleon’s progressive encoding strategy.
When approximations diverge from exact results, systems must assume the types of errors users find acceptable and find ways to convey the uncertainty. However, users struggle to interpret statistical uncertainty [38, 55, 42]. Further, accuracy guarantees over individual aggregates are not enough, because they may not preserve higher-level patterns that guide visual interpretation. BlinkDB [2] addresses these issues by exposing user-specified error or time budgets, but users find it hard to set these parameters. Distribution-oriented methods seek to preserve shapes or density profiles [21, 50], or to preserve properties of functions over the resulting visualization; for instance, Rahman et al. [63] preserves pairwise orderings in bar charts.
Unfortunately, perceptual error tolerances vary across people, tasks, and encodings and are not yet fully characterized. This makes property-driven approximations powerful but brittle. An open question is how to define perceptual models that are both general enough to apply across visualizations and precise enough to support optimization. Such models would allow systems to reason about approximation error in terms of perceived fidelity rather than statistical metrics. For instance, our work on perceptual functions (p-funks) [3] represents an early attempt to formalize these constraints. It specifies a just-noticeable difference (JND) model as function of mark properties (e.g., pixel height) so the system can guarantee that online approximations remain within perceptual bounds. This illustrates how interface-aware perceptual models can guide system behavior, but also highlights the need for more principled, general representations of visualization semantics.
@@ 2.2.3 A Deluge of Special Cases
Interface designers face a difficult and increasingly untenable position. To achieve interactive performance, they must choose layouts, encodings, and interaction patterns that align with a sprawling ecosystem of bespoke system optimizations—each tailored to a narrow performance sweet spot under specific interface or user assumptions. Simply understanding the subtle distinctions between dozens of approximation, caching, and scheduling strategies is challenging. Further deciding which combinations apply to a particular interface is beyond the reach of most developers.
@@ 2.3 Formal Representations & Design Independence
The brittle nature of special-case optimizations highlights the need for Design Independence: a principled way to represent interface semantics so the system, rather than the designer, can decide how to meet interactivity constraints. Design Independence would let interface designers rapidly iterate without being bottlenecked by checking with developers or architects about implementation feasibility [77], much as logical plans let query optimizers reason about execution without exposing physical details.
@@ 2.3.1 Physical Visualization Design
Physical Visualization Design (PVD) [14] is our approach towards Design Independence. Similar to physical database design (PDD) [17, 91, 25], which selects physical structures and layouts to accelerate a representative query workload, PVD selects data structures, caching strategies, and execution plans to meet the latency constraints of a visualization interface. The key idea is to provide a formal yet lightweight representation of the interface that captures the data flow, interactions, and latency requirements. This representation is input to an automated optimizer can analyze and synthesize a feasible client–server–DBMS execution strategy, and then instantiate the architecture.
@@
At first glance, PVD should not be possible. PDD is already considered intractable due to challenges of cost estimation and a combinatorial search space, and solutions only attempt to reduce expected workload latency. PVD appears to be strictly harder because it guarantees wall-clock latency for interactions. Surprisingly, we find that these millisecond-to-second latency constraints actually simplify the search because they eliminate the hardest cases in query optimization: joins with large or unpredictable fan-out. If a join cannot guarantee bounded output size [5], it cannot satisfy the latency constraints and is pruned. This dramatically reduces the search space, making PVD feasible despite its stronger guarantees.
@@ 2.3.2 An Instance of Physical Visualization Design
Our system, Jade [14], is the first instantiation of PVD that focuses on choosing and placing data structures for interactive visualization interfaces. Because modern visualizations rely on bespoke data structures (e.g., cubes, sketches, indexes), Jade is designed to support an extensible library of custom data structures.
@@
Figure 3(d) uses Cube.build() to compute a byte array that encodes a dense cube, the Cache() operator to reuse the cube when \$a or \$b change, and Cube.eval(2001, 2020) uses the cube to compute the vote distribution between 2001 and 2020. Data structures also estimate their cost and size based on statistics, and specify a match() method to identify subplans they can replace.
@@ 3.1 From Query Workloads to New Interactions
While popular, this library largely reflects what today’s systems can execute efficiently, rather than the full range of analytical operations users perform. For example, what if the user wishes to compare data slices, e.g., “the average cost over time for 75-year-olds against the overall population?” Such comparisons are commonplace in OLAP and cube-based analytics [29, 65] but have no counterpart in modern visualization systems because they require reasoning about the semantics and schemas of multiple derived views, not just editing a predicate in a single query.
@@
This gap motivated our View Composition Algebra work (VCA) [86], which treats views as first-class algebraic objects that can be composed, joined, merged, or differenced. Since views are the output of query pipelines, VCA operates over entire data-processing pipelines; this is in contrast to traditional interactions that modify SQL literals. Figure 8 demonstrates VCA in practice: the user drags the 2016 election results onto the 2020 results to compute their difference and visualize the result as a map—with no manual data wrangling. Figure 7 shows 4 illustrative VCA operators: (a) element-wise subtract each mark in one view with its match(es) in the other, (b) unioning the records in two views together, (c) computing the average of each date’s delays across all cities, and (d) implicitly joining the underlying tables to subtract the blue bar from every date.
Ideally, a user can freely compare any combination of data visualized in an interface. However, supporting such ad hoc comparisons in visual analysis tools is tricky because the specific data that will be visualized is often not known ahead of time and not all data is directly comparable. Thus, ad hoc comparisons require the system to dynamically assess whether a combination of views is valid to compose. From a data management perspective, this reduces to classic data integration problems: Are two views comparable? That is, do their schemas admit a semantically valid mapping (e.g., cost $\sim$ price)? If so, which tuples should be matched? This is entity resolution. Data integration remains an open problem, and thus, advances in schema matching, type inference, and entity resolution directly expand what comparison interactions can support.
Comparison is just one example of interaction techniques made possible by data-management concepts. Another is explanatory interactions, which surface predicates that explain surprising visual patterns [84]. But enabling explanations inside interactive systems required nearly a decade of systems innovations from our lab—to support fast lineage capture [60, 59, 58] and scalable lineage evaluation [52]—that combines with prior work on fast explanation discovery algorithms [1, 64, 28]. These systems-level innovations need to be in place before such interactions are feasible at the latency expected from modern interfaces.
@@ 3.2 From Query Workloads to New Interfaces
Our second project shows how the compositional structure in query workloads can be used to support automatic interface synthesis to go beyond today’s template-based approaches to interface design.
@@
We formalized this idea as the Data Interface Grammar (DIG) [15]. DIG models interfaces as grammars whose non-terminals correspond to interface choices and whose productions encode schema-aware SQL structure. DIG extends Parsing Expression Grammars (PEGs) [26] with SQL- and schema-specific annotations that constrain allowable expressions. A valid interface is a visual representation that expresses this grammar—user interactions resolve non-terminals to produce a concrete parse tree, which corresponds to a valid query.
@@
This representation offers several benefits. First, applications don’t assemble raw SQL strings and every query conforms to the grammar. Second, grammar rewrites offer a simple mechanism to refactor and explore alternative interface designs. Third, DIG aligns naturally with existing data engineering practices. For instance, DBT pipelines already structure analytics as DAGs of SQL views, which DIG can express and directly map to interactive dashboards. Fourth, DIG shares structural similarities with the plan-based representations used in PVD (Section 2.3), making it amenable to automatic optimization. Finally, standard techniques such as grammar induction can infer DIG grammars from query traces, enabling interface synthesis—an approach we explored in a line of work called Precision Interfaces [90, 16, 74, 13].
DIG is also useful when using large language models (LLMs). Natural language queries often contain unresolved ambiguities in intent, schema interpretation, or operator choice [36]. Rather than forcing the system to commit to a single query, the LLM can simply represent ambiguity as non-terminals and return an interface whose interactions navigate the space of ambiguities. This idea underlies NL2Interface [13], where LLMs generate a grammar expressing possible interpretations, and the user resolves them through interaction.